OpenStudy (anonymous):

Do not use spaces in your answer. If f(x) = (-x)^3, then f(-3) = _________.a^0

OpenStudy (solomonzelman):

\(f(x)=(-x)^3\) \(f(\color{red}{-3})=(-\color{red}{-3})^3~~~~\Rightarrow~~~~f(\color{red}{-3})=(3)^3=?\)

OpenStudy (solomonzelman):

3³ = 3 • 3 • 3

OpenStudy (anonymous):

27?

OpenStudy (solomonzelman):

Yup, very good!

OpenStudy (anonymous):

thanks

OpenStudy (solomonzelman):

Anytime;)